I have 12 coins that total the value of $2.25. If there are only quarters and dimes, how many of each kind are there
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

I have 12 coins that total the value of $2.25. If there are only quarters and dimes, how many of each kind are there.
----------
d + q = 12 coins
10d + 25q = 225 cents
----------------------------
2d + 2q = 24
2d + 5q = 45
----------------------
Subtract and solve for "q"::
3q = 21
q = 7 (# of quarters)
-------------------
d = 12-q = 12-7 = 5 (# of dimes)
----------------
Cheers,
Stan H.
-------------

Answer by addingup(3677)

q+d = 12, q = 12-d we'll use this value for q below
0.25q+0.10d = 2.25 substitute for q
0.25(12-d)+0.10d = 2.25
3-0.25d+0.10d = 2.25
-0.15d = -0.75 Divide both sides by 0.15, and remember -/- = +
d = 5 there are 5 dimes and 12-5 = 7 quarters
-------------------------------
check:
5*0.10 = 0.50
7*0.25 = 1.75
Total . .2.25 Correct
:
John